OpenStudy (anonymous):

Ryan earns $5 for each small dog he walks and $8 for each large dog he walks. Today he walked 9 dogs and made a total of $51. How many large dogs did Ryan walk?

OpenStudy (anonymous):

This is systems of equations

OpenStudy (anonymous):

s= small dog l= large dog

OpenStudy (xapproachesinfinity):

let x be the number of large dogs Ryan walks and y the number of small dogs x+y=9 8x+5y=51

OpenStudy (xapproachesinfinity):

you need to solve for x here

OpenStudy (anonymous):

s+l=9 5s=8l=51

OpenStudy (anonymous):

so I end up with 9x + 6y = 60 Where do I go from there?

OpenStudy (xapproachesinfinity):

No that's not how you solve the system

OpenStudy (xapproachesinfinity):

first write y in terms of x like this y=9-x then sub y into the second equation 8x+5(9-x)=51

OpenStudy (xapproachesinfinity):

solve for x to find the number of dogs you need

OpenStudy (anonymous):

Thank you!

OpenStudy (xapproachesinfinity):

welcome^_^

OpenStudy (xapproachesinfinity):

if you get x=2 then you got it correct!